OpenStudy (anonymous):

Integrate ∫(2 cos⁡x)/(1+sin^2 x)

hartnn (hartnn):

is it \(\huge \int \frac{2cos x}{1+sin^2 x}\) ?

hartnn (hartnn):

if yes, u can put t=sin x

OpenStudy (anonymous):

Yeah...i get it! Thank you v much :)

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

any more doubts / questions ?

OpenStudy (anonymous):

Umm..yes...integration is driving me crazy T.T What about this question... Integrate 1/ 1+sin^2 X use the same approach?

hartnn (hartnn):

hmm... no u divide numerator and denominator by cos^2 x and then put t=tan x if u see my tutorial, uget a hint on this..

hartnn (hartnn):

sec^x / (sec^2 x+1) t= tan x sec^2 x = 1+t^2 dt=.... ?

OpenStudy (anonymous):

Srry, i still cant get it...What i did is like this: 1/(1+sin^2 x)= (1/(cos^2 x))/((1+sin^2 x)/(cos^2 x)) = (sec^2 x)/(sec^2 x+tan^2 x)

hartnn (hartnn):

oh yes, my mistake now you can put t=tan x sec^2 x + tan^2 x= 1+tan^2x+tan^2x = 1+2t^2 and dt=.. ?

OpenStudy (anonymous):

So it bcome ∫〖1/(1+2t^2) dt ,am i right? Then i struck....

hartnn (hartnn):

right

hartnn (hartnn):

standard integral \(\int 1/(1+x^2)dx=tan^{-1}x\)

hartnn (hartnn):

here first write 2t^2 as \(\sqrt{2t}\)

hartnn (hartnn):

i mean \(\sqrt{2t}^2\)

hartnn (hartnn):

so u finally get \(tan^{-1}(\sqrt2 t)/\sqrt2+c\)

OpenStudy (anonymous):

This is what i did: ∫1/(1+(√2 t)^2) dt Let u=√2t dt=1/√2 t (1/√2) ∫1/(1+u^2) du (1/√2 ) tan^(-1)⁡u (1/√2) tan^(-1)⁡〖√2 t〗 and finally, i get this (1/√2 ) tan^(-1)⁡〖√2 tan⁡x 〗 Any mistake here?

hartnn (hartnn):

thats correct

OpenStudy (anonymous):

Alright.Thx a lot. U really help me a lot...:)

hartnn (hartnn):

welcome ^_^